Fermi National Laboratory
 

Accelerator Update for February 22, 2006 - February 24, 2006

 

Wednesday February 22

The day shift began with Operations monitoring store 4664 and the Recycler stash, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the store, with the Antiproton Source (Pbar) stacking, and with NuMI, MiniBooNE, and the Meson MTest and MCenter experiments taking beam.

Operations terminated store 4664 at 10:17 AM and then prepared D0 and CDF for an access into their collision halls.

Operations prepared the Recycler at 1:32 PM to stash antiprotons.

Operations put the Tevatron (TeV) through a dry squeeze at 2:09 PM.

The stash ended around 2:30 PM.

Operations put the TeV into shot setup around 2:57 PM.

Operations established store 4666 at 5:08 PM with an initial luminosity of 159.49E30. MCenter resumed taking beam.

MTest resumed taking beam at 5:49 PM.

At 6:06 PM, the TeV suffered a sector F4 quench.

The Recycler began preparing to stash around 6:39 PM. The stash ended. MTest and MCenter resumed taking beam at 7:39 PM. Pbar resumed stacking at 7:41 PM.

At 7:49 PM, Operations reported that the vacuum at the downstream end of sector F4 was degrading and the cryo system temperatures were rising.

At 7:49 PM, operators began preparing the TeV for a transfer hall access.

Operations prepared D0 for access at 8:28 PM.

At 8:48 PM, operators reported finding a blown relief valve on a TeV dipole.

Super Insulation was found on the floor. The quench was caused by a sloppy abort, but it wasn't until much later that a controls expert found the bad C200 module that had caused the abort.

At 10:38 PM, Operations reported that the Recycler had lost 18E10 of its stash. Recycler studiers stated that they understood what had happened.

Cryo system technicians reported that they were finished in the transfer hall and that TeV sector F4 was slowly warming up.

Thursday February 23

The midnight shift began with Operations monitoring the stash, with TeV sector F warming up, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

The Recycler began preparing to stash at 12:37 AM. The stash ended. Pbar resumed stacking at 1:36 AM. MTest and MCenter resumed taking beam at 1:44 AM.

Operations halted stacking at 1:50 AM.

The Neutron Therapy Facility (NTF) requested Linac beam to treat patients at 9:15 AM.

The Cryo System Run Coordinator reported at 10:48 AM that techs would be turning off all the TeV cold compressors and begin warming up sector A3.

The Recycler began preparing to stash at 8:07 PM. The stash ended. Pbar resumed stacking at 9:01 PM. MTest and MCenter resumed taking beam.

Pbar experts have completed their documentation of the Pbar beam. Stacking was halted at 11:19 PM.

Friday February 24

The midnight shift began with Operations monitoring the stash, with the TeV cryo systems warming up, with Pbar off, and with NuMI, MiniBooNE, MTest, and MCenter taking beam.

A Recycler expert began a study period at 2:32 AM.

The Recycler conducted a boomerang test with MI at 5:44 AM. The Recycler sent antiprotons to MI and MI sent them back at 100% efficiency. MI used the beam to calibrate their new BPM system.

Operations prepared D0 for access at 7:08 AM.

The Plans for Friday

The plans for today are to continue sending low intensity beam to NuMI, MiniBooNE, MTest, and MCenter until Sunday, 0000 hrs. Many of the machines will conduct studies. The twelve-week shutdown begins on Monday, February 27.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/28/2006    email Fermilab